HIGH RESOLUTION FAR INFRARED FOURIER TRANSFORM SPECTROSCOPY OF THE NH_{2} RADI-CAL.

\underline{M. A. MARTIN-DRUMEL}, O. PIRALI, D. BALCON, SOLEIL Synchrotron, AILES beamline, Saint-Aubin, France and Institut des Sciences Moleculaires d’Orsay, ISMO, CNRS, Universite Paris XI, Orsay, France; M. VERVLOET, SOLEIL Synchrotron, AILES beamline, Saint-Aubin, France.

First identified toward Sgr B2^{a}, the NH_{2} radical has recently been detected in the interstellar medium by the HIFI instrument on board of Herschel^{b}. Despite the fact that this radical has not been detected in brown dwarfs and exoplanets yet, it is already included in physical and chemical models of those environments^{c} (temperature higher than 2000 K expected in several objects). Its detection in those objects will depend on the existence of a reliable high temperature and high resolution spectroscopic database on the NH_{2} radical. The absorption spectrum of NH_{2} has been recorded between 15 and 700 cm^{−1} at the highest resolution available using the Bruker IFS125HR Fourier transform interferometer connected to the far infrared AILES beamline at SOLEIL (R=0.001 cm^{−1}). The radical was produced by an electrical discharge (DC) through a continuous flow of NH_{3} and He using the White-type discharge cell developped on the beamline (optical path: 24m). Thanks to the brilliance of the synchrotron radiation, more than 700 pure rotational transitions of NH_{2} have been identified with high N values (N_{max}=25) in its fundamental and first excited vibrational modes. By comparison to the previous FT spectroscopic study on that radical in the FIR spectral range^{d}, asymmetric splitting as well as fine and hyperfine structure have been resolved for several transitions.

aE. F. Van Dishoeck, D. J. Jansen, P. Schilke, T. G. Phillips The Astrophysical Journal 416, L83-L86 (1993)

bC. M. Persson, J. H. Black, J. Cernicharo et al. Astronomy and Astrophysics 521, L45 (2010)

cK. Lodders and B. Fegley, Jr Icarus 155, 393424 (2002)

dI. Morino and K. Kawaguchi Journal of Molecular Spectroscopy 182, 428-438 (1997)